OpenStudy (anonymous):

Find the four arithmetic means between 12 and -13. A. 6, -1, -3, -8 B. 6, 2, -3, -8 C. 7, -1, -3, -8 D. 7, 2, -3, -8

OpenStudy (luigi0210):

well arithmetic mean is adding/subtracting the same number..

OpenStudy (jim766):

first determine how far apart the numbers are 12 - (-13) = 25 then since you what 4 means, take 25 and divide by 4+1 or 5 25/5 = 5 -13 +5 -8 + 5 -3 +5 2 +5 7 +5 12 -13, -8, -3, 2, 7, 12 these are the 4 means

OpenStudy (luigi0210):

@Jim766 is right but none of the choices match up.. the closet is D..

OpenStudy (luigi0210):

wait nvm it is D

OpenStudy (luigi0210):

so the difference is 5